The hydrolysis of ethyl acetate in alkaline conditions yields acetate ion and ethanol. The rate

of the reaction increases 6.37 times its original value when the temperature is increased from

15°C to 45°C. Calculate the activation energy of this reaction.


1
Expert's answer
2021-02-09T04:36:49-0500

K1= 1

K2= 6.37

T1= 15°C = 288K

T2= 45°C = 318K

Ea=?

R= 8.314J/K.mol

Using the derived form of Arrhenius equation

lnK1/K2 = Ea/R x (1/T2 - 1/T1)

ln1/6.37 = Ea/8.314 x (1/318 - 1/288)

-1.85= Ea/8.314 x (-3.76 x 10-4)

Ea= -1.85x8.314/-3.76x10-4

Ea= 40907J = 41KJ